on 3 consecutive passes, a football team gain 7 yards, loses 31 yards, gains 24 yards. What number represents the total net net yardage? The total net yardage is ____ Yards.
Answer by drk(1908)

gain 7 yards means +7
loses 31 yards means -31
gains 24 yards means + 24
--
We have
+7 - 31 + 24
+7 - 31 = -24
-24 + 24 = 0.
--
What number represents the total net net yardage? 0
The total net yardage is 0 Yards.